# Going beyond density functional theory accuracy: Leveraging experimental data to refine pre-trained machine learning interatomic potentials

### v1

@@ Conclusions
In this work, we used a macroscopic experimental observable- EXAFS spectra- to refine the MLIP of two nuclear fuel systems (UO2 and UN) that were pre-trained using first-principles calculations (DFT+U/DFT). The choice of improving the MLIPs of nuclear fuel materials can serve a larger purpose, by helping to accelerate fuel qualification, which aims to explore new nuclear fuel candidates while reducing the number of nuclear integral tests required. The accurate predictions by these proposed models can be used to build data science guided mechanistic models that aid in the application of multi-physics reactor modeling.
Although the pre-trained MLIPs show good agreement with the respective DFT/DFT+U energies and forces, the agreement with certain experimental properties, e.g., lattice parameter, is limited by the accuracy of the DFT/DFT+U training data. Because EXAFS spectra strongly depend on the interatomic distances, coordination number, and atomic arrangement around an absorbing atom, improving the MLIP to match the target EXAFS spectra particularly improves structural properties. Other properties, such as diffusion coefficient and defect energies remain relatively unchanged, demonstrating that the re-trained MLIP does not βforgetβ what was learned from the DFT/DFT+U dataset. Both, UO2 and UN refined MLIPs are tested for properties such as 0Β K elastic constants, temperature-dependent lattice parameters, heat capacity, bulk modulus, and defect energies. Additionally, the refined UO2 MLIP is tested for the activation energy of oxygen diffusion and phonon dispersion spectra. In general, both MLIPs retain the robustness from pre-training to DFT/DFT+U data while improving the agreement with the experimental EXAFS spectra and additional properties.
@@
This work advances the limits of training MLIPs only to quantum mechanical data by refining a pre-trained MLIP to match experimental data. Future work will investigate simultaneous training to both data sources and/or multiple experimental sources. The current refinement method is limited to properties that can be computed for every atom for a given snapshot in a simulation, such as EXAFS spectra. Future work will investigate the application of this approach to other properties, such as radial distribution functions and pressure-volume equations of state.
@@
All EXAFS spectra calculations are performed using the ab initio real-space multiple-scattering code FEFF8.5[60, 39, 61]. The scattering paths only within the radius 6 Γ of the absorbing atom for both UO2 and UN are considered. To develop the target synthetic EXAFS spectra, FEFF calculations are performed for U and O atom in a supercell UO2 system and for U and N atoms in a UN supercell. EXAFS corresponding to the K-edge absorption is calculated for O and N, while the L3-edge spectra is calculated for U in both materials. Self-consistent field (SCF) method for calculations improves the agreement of the synthetic EXAFS spectra with the experimental spectra. However, performing SCF is computationally expensive, hence we did not implement SCF for any target spectrum or during the refinement process. SCF was only employed when computing a spectrum for comparison with the experimental EXAFS spectrum. The FEFF parameters used to generate the target synthetic EXAFS spectra using perfect crystal, and the EXAFS spectra of the MD trajectories during the refinement process are kept consistent. FEFF input headers are provided in Supporting Information.
@@
The MLIPs for UO2 and UN pre-trained to DFT+U and DFT datasets, respectively, were refined by transfer learning to match the synthetic target EXAFS spectra. Ideally, these target EXAFS spectra would be obtained from experiments. Unfortunately, experimental EXAFS spectra are available only for the uranium atom in UO2 and UN. For this reason, we generated synthetic target EXAFS spectra for both elements (U and O in UO2 and U and N in UN) using an ideal stoichiometric perfect crystal. To improve the agreement between synthetic spectra and experimental spectra at finite temperature, the perfect crystal is constructed using the experimental lattice parameter at 75 K for UO2 and at 50 K for UN.
@@
MLIP refinement is a three-step process. First, molecular dynamics simulations are performed with the MLIP undergoing refinement, i.e., the original unrefined MLIP for the first refinement generation and preliminary refined MLIPs for subsequent generations. Second, the atomic energy shifts () are optimized to improve agreement between the reweighted spectra and the target EXAFS spectra (by minimizing the loss function equationΒ 5). Third, the MLIP is re-trained using transfer learning to the shifted atomic energies.
@@
MD simulations for EXAFS calculations were performed in the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS)[62] using a 4 4 4 supercell of UO2 (768 atoms) and 5 5 5 supercell of UN (1000 atoms). We first relax the simulation boxes and then perform isobaric-isothermal simulations (NPT) at constant pressure P=0 bar and temperature T = 75 K for UO2 and T = 50 K for UN using a NosΓ©-Hoover thermostat with damping at every 100 timesteps and barostat with damping at every 1000 timesteps, for 16 ps with a timestep of 0.2 fs.[63] After equilibration, eight independent simulations are performed, each with random initial velocity, for 48 ps with a timestep of 0.2 fs. Atomic configurations and the corresponding atomic energies (E0) are saved for three snapshots from each of the eight MD runs after every 16 ps. Because each snapshot contains either 768 (UO2) or 1000 (UN) atoms, the total number of atomic energies in the refinement dataset is 18432 for UO2 and 24000 for UN.
@@
EXAFS spectrum is calculated for each atom in the refinement dataset using the FEFF8.5 code[60, 39, 61]. The spectrum is computed only for the wavevector range k = 2.05 to 14.95, with an interval of 0.05. Hence, each spectrum is an array of size 259. Two loss functions (EquationΒ 5) are minimized separately for U atoms and O atoms in UO2 to obtain the shifted atomic energies. In EquationΒ 5, the size of vector is equal to the length of a single spectrum. The number of rows in and corresponds to the length of a single spectrum, and the number of columns is equal to the total number of spectra. For materials with a single element, the total number of spectra is equal to the total number of atoms in the refinement dataset, and a single loss function is defined. Both systems studied here have two element types, each with a different target EXAFS spectrum, for which two loss functions are defined. Therefore, the number of columns in in each loss function represents the number of atoms of the same element type in the refinement dataset. The length of the re-weighting vector () is also equal to the number of same-element atoms in the refinement dataset. The simulated spectra matrix () contains data from all selected snapshots and has dimensions = 259 6144 for the loss function corresponding to U, and dimensions = 259 12288 for the loss function corresponding to O in UO2. Similarly, two loss functions are minimized for U atoms and N atoms in UN, where both U and N have () dimensions = 259 12000.
@@
Our refinement process utilizes transfer learning to avoid overfitting to the refinement dataset and so that the refined MLIP retains some of the information learned from the DFT/DFT+U training dataset. Our transfer learning approach consists of three key aspects. First, the refined MLIP is initialized with the pre-trained MLIP parameters. Second, retraining is performed with a small learning rate (10-4) for a maximum of five total epochs. Third, certain MLIP parameters are not fine-tuned during retraining. Specifically, for UO2, the atomic layers (0 to 3) are kept frozen. For UN, however, no layers were frozen because the pre-trained MLIP required smaller changes and, thus, the refinement process was sufficiently stable without freezing any layers.
@@
All MD simulations of temperature-dependent properties reported in this work are performed with the LAMMPS software, unless stated otherwise.[62] We employed a larger 12 12 12 supercell (20736 atoms in UO2, 13824 atoms in UN) for better statistics and to avoid finite-size effects. The MD simulations are equilibrated using the isobaric-isothermal ensemble (NPT) with constant pressure of 0 bar and constant temperature in the range between 225 K and 2575 K (for UN) and from 225 K to 3075 K (for UO2) using 25 K increments. We use a NosΓ©-Hoover thermostat with damping factor of 0.1 ps and barostat with anisotropic pressure control and a damping factor of 0.5 ps[63]. Timestep was set to 2 fs and simulations were performed for 20 ps. For extracting the properties of interest, we extract the enthalpy (H) and lattice parameter by averaging over the last 8 ps of the simulation. The partial derivative of enthalpy with respect to temperature at constant pressure is used to calculate the heat capacity as:
@@
Bulk modulus is calculated every 100 K in a temperature range of 300 K to 2500 K for UN, and from 300 K o 3000 K for UO2. The nominal volume is calculated by averaging over the last 4 ps of the NPT MD run at the respective temperature. Subsequently, these equilibrated structures are subject to compression and expansion by 1% and 2% each, with respect to the nominal volume . The compression and expansion MD simulations are performed using the NVT ensemble (constant volume and temperature). Each NVT simulation is also held for 20 ps using a timestep of 2 fs. The pressure for a given compression or expansion is averaged over the last 4 ps of each run. The bulk modulus is defined as
where the partial derivative of pressure with respect to volume is evaluated at the nominal volume, . To estimate this partial derivative, the slope of P with respect to V at a given T is calculated from the linear fit, using five data points - two from compression, two from expansion, and one from zero pressure - at the same simulation temperature.
The diffusion coefficient of oxygen atoms in UO2 is calculated using NernstβEinstein equation, as shown below:
where is the mean-square-displacement (MSD) and is time. This derivative is calculated as the linear slope of MSD with respect to time (see Figure S13). To calculate MSD, the UO2 structures are equilibrated in the NPT ensemble for 20 ps from 2100 K to 3000 K with 100 K increments. We then perform a constant energy simulation using the NVE ensemble for 1 ns with a timestep of 2 fs. Transport properties are best computed with the NVE ensemble so that the dynamics are not impacted by the barostat or thermostat[64]. MSD values are recorded every 200 fs from the generated trajectory.
The activation energy () is obtained from the temperature-dependent diffusion coefficient using linear regression to the logarithmic form of the Arrhenius equation:
@@
The point defect energies are evaluated for a 4 4 4 UO2 supercell (768 atoms) and a 5 5 5 UN supercell (1000 atoms). For the Schottky defect (SD) in UO2, a cation U4+ and two anions O2- are removed from the crystal structure. In the case of UN, a cation U3+ and an anion N3- are removed from the crystal structure. The SD defect energy is calculated as:
@@
The Frenkel pair defect (FP) is defined separately for a cation and an anion. In either case, a charged species is removed from its lattice position and placed at an interstitial site, thus retaining a charge-neutral system. For both UO2 and UN, the interstitial position to place the cation / anion is selected randomly. Since is equal to in the case of FP defect, the FP defect energy is calculated as:
The phonon dispersion spectrum at 0 K is computed for a 10 10 10 UO2 supercell using the finite difference method as implemented in the Atomic Simulation Environment (ASE) package.[65]

### v2

@@ Conclusions
In this work, we used a macroscopic experimental observable–EXAFS spectra–to refine the MLIP of two nuclear fuel materials, UO2 and UN, that were pre-trained using first-principles calculations (DFT+U and DFT, respectively). The choice of improving MLIPs developed for nuclear fuel materials can be useful for accurate atomic-level predictions that can improve mechanistic and multi-physics reactor models used in accelerated fuel qualification–a methodology that aims to reduce the number of nuclear integral tests required in order to qualify a new fuel. Furthermore, this work presents a real-world application typical of computational materials modeling where the DFT+U/DFT calculations are qualitatively reliable for pre-training purposes.
Although the pre-trained MLIPs show good agreement with the respective DFT/DFT+U energies and forces, the agreement with certain experimental properties, e.g., lattice parameter, is limited by the accuracy of the DFT/DFT+U training data. Because EXAFS spectra strongly depend on the interatomic distances, coordination number, and atomic arrangement around an absorbing atom, improving the MLIP to match the target EXAFS spectra particularly improves structural properties. Other properties, such as diffusion coefficient and defect energies remain relatively unchanged, demonstrating that the re-trained MLIP does not “forget” the original learning from the DFT/DFT+U dataset. Both, UO2 and UN refined MLIPs are tested for properties such as elastic constants, temperature-dependent lattice parameters, heat capacity, bulk modulus, and defect energies. Additionally, the refined UO2 MLIP is tested for the activation energy of oxygen diffusion and phonon dispersion spectra.
In general, both MLIPs retain the robustness from pre-training to DFT/DFT+U data while improving the agreement with experimental data. In particular, the predictions for the UO2 refined MLIP improve significantly for the lattice parameter and bulk modulus over the entire temperature range, while marked improvement is also observed for heat capacity at high temperatures, thermal expansion coefficient at high temperatures, elastic constants at 300 K, and the experimental EXAFS spectra at both 300 K and near the refinement temperature. For UN, the improvement is most pronounced for the lattice parameter over the entire temperature range, while improvement is also observed for heat capacity at high temperatures, elastic constants at 300 K, and the experimental EXAFS spectra at both 300 K and near the refinement temperature.
@@
This work advances the limits of training MLIPs only to quantum mechanical data by refining a pre-trained MLIP to match experimental data. Future work will investigate simultaneous training to both data sources and/or multiple experimental sources. The current refinement method is limited to properties that can be computed for every atom for a given snapshot in a simulation, such as EXAFS and RDF spectra. The procedures pioneered in this paper have a wide range of applicability, including to complex microstructures, e.g., interfaces and grain boundaries, as well as to molecular systems. Future work will investigate the application of this approach to other properties, such as angular distribution functions and pressure-volume equations of state.
@@
All EXAFS spectra calculations are performed using the ab initio real-space multiple-scattering code FEFF8.5 [2, 48, 49]. The scattering paths only within the radius 6 Å of the absorbing atom for both UO2 and UN are considered. To develop the target synthetic EXAFS spectra, FEFF calculations are performed for uranium and oxygen atom in a supercell of the UO2 crystal (768 atoms) and, for uranium and nitrogen atoms in a UN supercell (1000 atoms). EXAFS corresponding to the K-edge absorption is calculated for O and N, while the L3-edge spectra is calculated for U in both materials. The self-consistent field (SCF) method for calculations improves the agreement of the synthetic EXAFS spectra with the experimental spectra. However, performing SCF is computationally expensive, hence we did not implement SCF for any target spectrum or during the refinement process. SCF was only employed when computing a spectrum for comparison with the experimental EXAFS spectrum. The FEFF parameters used to generate the target synthetic EXAFS spectra using perfect crystal, and the EXAFS spectra of the MD trajectories during the refinement process are kept consistent. FEFF input headers are provided in Supporting Information.
@@
The MLIPs for UO2 and UN pre-trained to DFT+U and DFT datasets, respectively, were refined by transfer learning to match the synthetic target EXAFS spectra. Ideally, these target EXAFS spectra would be obtained from experiments. Unfortunately, experimental EXAFS spectra are available only for the uranium atoms in UO2 and UN. For this reason, we generated synthetic target EXAFS spectra for both elements (U and O in UO2, and U and N in UN) using an ideal stoichiometric perfect crystal. To improve the agreement between synthetic spectra and experimental spectra at finite temperature, the perfect crystal is constructed using the experimental lattice parameter at 75 K for UO2 and at 50 K for UN.
@@
MLIP refinement is a three-step process. First, molecular dynamics simulations are performed with the MLIP undergoing refinement, i.e., the original unrefined MLIP for the first refinement generation and preliminary refined MLIPs for subsequent generations. Second, the atomic energy shifts () are optimized to improve agreement between the reweighted spectra and the target EXAFS spectra (by minimizing the loss function equation 5). Third, the MLIP is re-trained using transfer learning to the shifted atomic energies.
@@
MD simulations were performed to generate configurations for EXAFS spectra calculations in the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [61] using a 4 4 4 supercell of UO2 (768 atoms) and 5 5 5 supercell of UN (1000 atoms). First, the atomic positions are relaxed to minimum energy configuration. Then, initial velocities are assigned based on a Maxwell-Boltzmann distribution. Next, isobaric-isothermal (NPT) simulations are performed for 16 ps with a timestep of 0.2 fs. A Nosé-Hoover barostat maintains a constant pressure of 0 bar with damping at every 1000 timesteps (0.2 ps). A Nosé-Hoover thermostat maintains constant temperature of 75 K for UO2 and 50 K for UN, with damping at every 100 timesteps (0.02 ps) [14]. After equilibration, eight independent production simulations in the NPT ensemble are performed for 48 ps. Three atomic configurations (snapshots) and the corresponding atomic energies (E0) are saved every 16 ps for each of the eight MD runs. Because each snapshot contains either 768 (UO2) or 1000 (UN) atoms, the total number of atomic energies in the refinement dataset is 18432 for UO2 and 24000 for UN.
@@
EXAFS spectra are calculated for each atom in the refinement dataset using the FEFF8.5 code [2, 48, 49]. The spectra are computed only for the wavevector range k = 2.05 to 14.95, with an interval of 0.05. Hence, each spectrum is an array of size 259, consistent with the dimensions of the target synthetic spectra and the experimental spectra. Two loss functions (Equation 5) are minimized separately for U atoms and O atoms in UO2 to obtain the shifted atomic energies. In Equation 5, the size of vector is equal to the length of a single spectrum. The number of rows in and corresponds to the length of a single spectrum, and the number of columns is equal to the total number of spectra. For materials with a single element, the total number of spectra is equal to the total number of atoms in the refinement dataset, and a single loss function is defined. Both systems studied here have two element types, each with a different target EXAFS spectrum, for which two loss functions are defined. Therefore, the number of columns in in each loss function represents the number of atoms of the same element type in the refinement dataset. The length of the re-weighting vector () is also equal to the number of same-element atoms in the refinement dataset. The simulated spectra matrix () contains data from all selected snapshots and has dimensions = 259 6144 for the loss function corresponding to U, and dimensions = 259 12288 for the loss function corresponding to O in UO2. Similarly, two loss functions are minimized for U atoms and N atoms in UN, where both U and N have () dimensions = 259 12000.
@@
Our refinement process utilizes transfer learning to avoid overfitting to the refinement dataset and so that the refined MLIP retains some of the information learned from the DFT/DFT+U training dataset. This approach consists of three key aspects. First, the refined MLIP is initialized with the pre-trained MLIP parameters. Second, retraining is performed with a small learning rate (10-4) for a maximum of five total epochs. Third, certain MLIP parameters are not fine-tuned during retraining. Specifically, for UO2, the atomic layers (0 to 3) are kept frozen. For UN, however, no layers were frozen because the pre-trained MLIP required smaller changes and contains nearly five times fewer model parameters than the UO2 MLIP. Thus, the refinement process for UN was sufficiently stable without freezing any layers.
We tested the efficiency of our refinement process by varying the amount of atomic energies in the refinement dataset by an order of magnitude. Specifically, we tested simulating systems between 324 and 2592 atoms of UO2 and saving between 2 and 20 snapshots per MD simulation. In all cases, the refinement process converges to the experimental lattice parameter at 300 K after approximately six generations of MLIP refinement. However, the larger datasets (more snapshots and/or more atoms) resulted in significantly more overfitting (without freezing any MLIP parameters), as evidenced by extremely poor prediction of lattice parameters at higher temperatures, bulk modulus at low temperatures, and heat capacity over the entire temperature range. For this reason, we recommend simulating relatively small systems and saving only a few snapshots per simulation.
Most of the refinement process is embarrassingly parallel and, thus, scales linearly with respect to the number of MD simulations or the number of saved snapshots per trajectory. Specifically, each MD simulation is performed in parallel on a single GPU, while FEFF calculations are performed in parallel on a single CPU. Thus, the real time to solution for MD and FEFF remains constant when increasing the number of GPUs and CPUs with respect to the size of the refinement dataset. However, our current code does not allow for multi-GPU training of HIP-NN. Fortunately, the retraining step is relatively fast as only a few parameter updates are performed each iteration. Future implementation of multi-GPU training will improve scalability for the MLIP refinement as well, resulting in an overall process with completely linear scaling.
@@
All MD simulations of temperature-dependent properties reported in this work are performed with the LAMMPS software, unless stated otherwise [61]. We employed a larger supercell for both systems (20736 atoms in UO2, 13824 atoms in UN) to avoid finite-size effects and provide better statistics. A set of MD simulations are performed every 25 K for 20 ps using the isobaric-isothermal ensemble (NPT) with constant pressure (0 bar) and constant temperature in a range between 225 K to 3075 K for UO2 and 225 K and 2575 K for UN. Timestep was set to 2 fs. The Nosé-Hoover thermostat was used with damping factor of 0.1 ps and barostat with anisotropic pressure control and a damping factor of 0.5 ps [14] for all simulations. The lattice parameters are extracted from simulation by averaging the last 8 ps of simulation. To compute heat capacity , the enthalpy (H) was obtained by averaging the last 8 ps of each run and used to calculate the partial derivative of enthalpy with respect to temperature at constant pressure as:
@@
Bulk modulus is calculated every 100 K between 300 K and 3000 K for UO2, and from 300 K to 2500 K for UN. The nominal volume is calculated by averaging the last 4 ps of the NPT MD run at each respective temperature. Subsequently, these equilibrated structures are subject to compression and expansion by 1% and 2% each, with respect to the nominal volume . The compression and expansion MD simulations are performed using the NVT ensemble (constant volume and temperature). Each NVT simulation is also held for 20 ps using a timestep of 2 fs. The pressure for a given compression or expansion is averaged over the last 4 ps of each run. The bulk modulus is calculated as:
where the partial derivative of pressure with respect to volume is evaluated at the nominal volume . To estimate this partial derivative, the slope of P with respect to V at a given T is calculated from a linear fit using five data points–two from compression, two from expansion, and one from zero pressure–at the same simulation temperature.
To calculate the near room temperature elastic constants, a 4 4 4 supercell of UO2 (768 atoms), and a 5 5 5 supercell of UN (1000 atoms) is equilibrated in the NPT ensemble at 296 K for UO2 and 290 K for UN for 500 ps, with a timestep of 1 fs. Nosé-hoover thermostat and barostat was used to held the temperature and pressure constant. The born/matrix method was used to calculate the tensor matrix [69].
The diffusion coefficients for oxygen in UO2 is calculated using the Nernst–Einstein equation:
where is the mean-square-displacement (MSD) and is time. This derivative is calculated as the linear slope of MSD with respect to time (see Figure S16). To calculate MSD from MD, the UO2 structures were equilibrated in the NPT ensemble for 20 ps from 2100 K to 3000 K with 100 K increments. We then perform a constant energy simulation using the NVE ensemble for 1 ns with a timestep of 2 fs. Transport properties are best computed with the NVE ensemble so that the dynamics are not impacted by the barostat or thermostat [38]. MSD values are recorded every 200 fs from the generated trajectory.
The activation energy is obtained from a linear fit of the diffusion coefficient as a function of temperature following an Arrhenius equation:
@@
Point defect energies are evaluated for a 4 4 4 UO2 supercell (768 atoms) and a 5 5 5 UN supercell (1000 atoms). For the Schottky defect (SD) in UO2, a cation U4+ and two anions O2- are removed from the crystal structure. In the case of UN, a cation U3+ and an anion N3- are removed from the crystal structure. The SD defect energy is calculated as:
@@
Frenkel pair defect energies (FP) are defined separately for a cation and an anion. In either case, a charged species is removed from its lattice position and placed at an interstitial site, thus retaining a charge-neutral system. For both UO2 and UN, the interstitial position to place the cation/anion is selected randomly. Since is equal to in the case of FP defect, the FP defect energy is calculated as:
The UO2 phonon dispersion band at 0 K is computed for a supercell using the finite difference method as implemented in the Atomic Simulation Environment (ASE) package [24].
Appendix
Aluminum Example
Our refinement process to experimental data is not limited to only EXAFS data, but can be applied to any experimental observable where the value is an average of responses from individual atoms. To highlight this fact, early in the development of the methodology presented in this manuscript, we implemented our refinement approach to compare with the refinement procedure of Matin et. al. [39] Specifically, we refined a DFT-level pre-trained MLIP for aluminum using experimental liquid RDF spectra. Similar to Matin et al., our refined MLIP predicts improved RDF spectra as well as improved diffusion coefficients. However, our refinement process based on modified reweighting has two key advantages over the procedure in Matin et al. First, the MLIP parameters are modified directly themselves, rather than introducing an additional pairwise potential. Second, our methodology is capable of refining to multiple temperatures simultaneously. In this section, we present results for refining the aluminum MLIP.
The unrefined MLIP is initially trained to an existing DFT-level aluminum dataset.[56] For this example, the MLIP is also based on the HIP-NN architecture with similar hyperparameters to the UO2 and UN MLIPs (see Supporting Information). Next, MD simulations are performed at two different temperatures for which the experimental RDFs are available, namely 1023 K and 1323 K.[67] The simulations were equilibrated for 200 ps using a 0.5 fs timestep. After equilibration, production MD simulations are run for 250 ps in the isobaric-isothermal (NPT) ensemble. During production, a binned RDF for every atom is saved every 5 ps. These binned atomic RDFs (ARDFs) extend to 10.0Å and contain 200 evenly spaced bins. The binned ARDFs are simply a count of the number of atoms around the central atom in a given radius , i.e., between r and r+ r. To improve statistical averaging and independent sampling, four MD simulations were run at each temperature. Each simulation contained 2048 atoms, resulting in a total of 819,200 binned atomic RDFs. ARDFs computed for a single atom from a single snapshot are intrinsically noisy. However, the RDF remains smooth and can be recovered by averaging over all snapshots and all atoms.
where is the number of binned ARDFs, is the distance at which the ARDF is evaluated, and is the instantaneous number density of the simulation at the corresponding snapshot.
Utilizing this reconstruction rule, the individual ARDFs were employed in Eq. 6 as and the same reweighting procedure was employed to determine the values. This procedure is performed in parallel at each temperature independently. The values are then added to the HIP-NN atomic energies to obtain the shifted atomic energies for each atomic configuration from the MD simulations, exactly as described in the main text. HIP-NN was then trained to the shifted atomic energies for 5 epochs with a learning rate of 1e-4. Once the next generation of MLIP was trained, new MD simulations were performed and the whole refinement process was repeated a total of nine times.
As shown in Figure 8, the refined MLIP predicts a total RDF in good agreement with the experimental data. Further, the simulated diffusion rates show significant improvement to the experimental values. The intermediate MLIPs show systematic improvement in the RDF and diffusion coefficient at the higher temperature with respect to generations of refinement. However, the non-monotonic improvement in the RDF and diffusion coefficient at the lower temperature is due to the refined MLIP occasionally raising the melting point near or above 1023 K, causing a precipitous drop in the diffusivity. However, this increase in the MLIP melting point is immediately corrected with the next generation of experimental refinement. Overall, this further example shows that the refinement procedure is generalizable to different experimental observables and can be applied to multiple experimental measurements (e.g., temperatures) at once.
Declarations